# Faster Gradient Methods for Highly-smooth Stochastic Bilevel Optimization

### v1

@@ 3 The F2SA-$p$ Method
In this section, we introduce our improved method for stochastic first-order bilevel optimization. In Section 3.1, we establish the relationship between fully first-order hyper-gradient approximation and finite difference schemes. In Section 3.2, we design our algorithm inspired by the $p$th-order finite difference method. In Section 3.3, we present the theoretical analysis of our algorithm and show that it achieves a $\tilde{\mathcal{O}}(p\epsilon^{-4-2/p})$ upper bound on SFO calls.
@@ 3.1 Hyper-Gradient Approximation via Finite Difference
The core idea of F2SA [28] is to solve the reformulated penalty problem (3) and use the gradient of the penalty function to approximate the true hyper-gradient. In this subsection, we revisit their hyper-gradient approximation and make connections to the finite difference method, which further motivates us to design better algorithms by using better finite difference formulas.
We start from the following relationship of the difference between perturbed lower-level problems and the original problem. We let
@@
where $g_{\nu}$ is the perturbed lower-lever problem with ${\bm{y}}_{\nu}^{*}({\bm{x}})$ and $\ell_{\nu}({\bm{x}})$ being its optimal solution and optimal value, respectively. Since the constraint ${\bm{y}}=\arg\min_{{\bm{z}}\in{\mathbb{R}}^{d_{z}}}g({\bm{x}},{\bm{z}})$ is equivalent to requiring $g({\bm{x}},{\bm{y}})\leq\min_{{\bm{z}}\in{\mathbb{R}}^{d_{y}}}g({\bm{x}},{\bm{z}})$, it can be shown [10, Lemma B.3] that $\frac{\partial}{\partial\nu}\ell_{\nu}({\bm{x}})|_{\nu=0}=\varphi({\bm{x}})$ holds under Assumption 2.3 and 2.4. Furthermore, Kwon et al. [28] showed that the partial derivatives with respect to $\bm{x}$} and $\nu$ are commutable, which leads to
@@
Let $\nu=1/\lambda$ in Eq. (3). Then the fully first-order hyper-gradient estimator [28, 10] is exactly using forward difference to approximate $\nabla\varphi({\bm{x}})$, that is,
@@
However, the forward difference is not the only way to approximate a derivative. Essentially, it falls into a general class of $p$th-order finite difference [4] that can guarantee an ${\mathcal{O}}(\nu^{p})$ approximation error. We restate this known result in the following lemma and provide a self-contained proof in Appendix B for completeness.
@@
Assume the unitary function $\psi:{\mathbb{R}}\rightarrow{\mathbb{R}}$ has $C$-Lipschitz continuous $p$th-order derivative. If $p$ is even, there exist $p$th-order central difference coefficients $\{\alpha_{j}\}_{j=-p/2}^{p/2}$ such that
$\left|\frac{1}{\nu}\sum_{j=-p/2}^{p/2}\alpha_{j}\psi(j\nu)-\psi^{\prime}(0)\right|={\mathcal{O}}(C\nu^{p}),$
where $\alpha_{0}=0$ and $\alpha_{j}=\alpha_{-j}$ for all $j=1,\cdots,p/2$. If $p$ is odd, there exist $p$th-order forward difference coefficients $\{\beta_{j}\}_{j=0}^{p}$ such that
$\left|\frac{1}{\nu}\sum_{j=-p/2}^{p/2}\beta_{j}\psi(j\nu)-\psi^{\prime}(0)\right|={\mathcal{O}}(C\nu^{p}),$
When $p=1$, we have $\beta_{0}=-1$, $\beta_{1}=1$, and we obtain the forward difference estimator $\psi(\nu)-\psi(0)/\nu$; When $p=2$ we have $\alpha_{-1}=-1/2,\alpha_{1}=1/2$ and we obtain the central difference estimator $(\psi(\nu)-\psi(-\nu))/(2\nu)$. Lemma 3.1 tells us that in general we can always construct a finite difference estimator ${\mathcal{O}}(\nu^{p})$ error with $p$ points with for even $p$ or $p+1$ points for odd $p$ under the given smoothness conditions. Inspired by Lemma 3.1 and Eq. (8) that $\frac{\partial^{2}}{\partial\nu\partial{\bm{x}}}\ell_{\nu}({\bm{x}})|_{\nu=0}=\nabla\varphi({\bm{x}})$, we propose a fully first-order estimator via a linear combination of $\frac{\partial}{\partial{\bm{x}}}\ell_{j\nu}({\bm{x}})$ to achieve ${\mathcal{O}}(\nu^{p})$ approximation error to $\nabla\varphi({\bm{x}})$ given that $\frac{\partial^{p+1}}{\partial\nu^{p}\partial{\bm{x}}}\ell_{\nu}({\bm{x}})$ is Lipschitz continuous in $\nu$. It further leads to Algorithm 1 that will be formally introduced in the next subsection.
Remark 3.1.
A subtlety to use Lemma 3.1 for hyper-gradient estimation is that it only applies to a unitary function while $\frac{\partial}{\partial{\bm{x}}}\ell_{\nu}({\bm{x}})$ is a vector-valued function in $\nu$. However, the approximation error still holds for the whole vector under the Euclidean norm if we apply the lemma on each dimension and note that the finite difference coefficients are the same for all dimensions.
@@ 3.2 The Proposed Algorithm
3:  Sample random i.i.d indexes $\{(\xi^{y}_{j},\zeta^{y}_{j})\}_{j=-p/2}^{p/2}$ and $\{({\xi_{i}^{x}},{\zeta_{i}^{x}})\}_{i=1}^{S}.$
4:  for $j=-p/2,-p/2+1,\cdots,p/2$
5:   ${\bm{y}}_{t}^{j,0}={\bm{y}}_{t}^{j}$
6:   for $k=0,1,\cdots,K-1$
@@
10:  end for
11:  Let $\{\alpha_{j}\}_{j=-p/2}^{p/2}$ be the $p$th-order central coefficients defined in Lemma 3.1.
12:  $\Phi_{t}=\frac{1}{S}\sum_{i=1}^{S}\sum_{j=-p/2}^{p/2}\alpha_{j}\left(jF_{x}({\bm{x}}_{t},{\bm{y}}_{t+1}^{j};\xi^{x}_{i})+\dfrac{G_{x}({\bm{x}}_{t},{\bm{y}}_{t+1}^{j};\zeta^{x}_{i})}{\nu}\right)$
13:  ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta_{x}{\Phi_{t}}/{\|\Phi_{t}\|}$
14:end for
Due to space limitations, we only present Algorithm 1 designed for even $p$ in the main text. The algorithm for odd $p$ can be designed similarly, and we defer the concrete algorithm to Appendix D. Our Algorithm 1 follows the double-loop structure of F2SA [10, 29] and changes the hyper-gradient estimator to the one introduced in the previous section. Now, we give a more detailed introduction to the procedures of the two loops of F2SA-$p$.
@@
Compared to [10, 28], the only modification we make to the outer loop is to change the gradient step to a normalized gradient step. The normalization can control the change of ${\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})$ and make the analysis of inner loops easier. We believe that all our theoretical guarantees also hold for the standard gradient step via a more involved analysis.
@@ 3.3 Complexity Analysis
This section contains the complexity analysis of Algorithm 1. We first derive the following lemma from the high-dimensional Faà di Bruno formula [32].
@@
Let $\nu\in(0,1/(2\kappa)]$. For any instance in ${\mathcal{F}}^{\text{nc-sc}}(L_{0},\cdots,L_{p+1},\mu,\Delta)$ as per Definition 2.2, $\frac{\partial^{p+1}}{\partial\nu^{p}\partial{\bm{x}}}\ell_{\nu}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{2p+1}\bar{L})$-Lipschitz continuous in $\nu$.
Our result generalizes the prior result for $p=1$ [28] to any $p\in{\mathbb{N}}_{+}$ and also tightens the prior bounds for $p=2$ [10] as we remark in the following.
@@
Note that the variables $\bm{x}$} and $\nu$ play equal roles in our analysis. Therefore, our result in $p=2$ essentially implies that $\frac{\partial^{3}}{\partial\nu\partial{\bm{x}}^{2}}\ell_{\nu}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{5}\bar{L})$-Lipschitz continuous in $\nu$ around zero, which tightens the ${\mathcal{O}}(\kappa^{6}\bar{L})$ bound of Hessian convergence in [10, Lemma 5.1a] and is of independent interest. The main insight is to avoid the direct calculation of $\nabla^{2}\varphi({\bm{x}})=\frac{\partial^{3}}{\partial\nu\partial{\bm{x}}^{2}}\ell_{\nu}({\bm{x}})|_{\nu=0}$ which involves third-order derivatives and makes the analysis more complex, but instead always to analyze it through the limiting point $\lim_{\nu\rightarrow 0_{+}}\frac{\partial^{3}}{\partial\nu\partial{\bm{x}}^{2}}\ell_{\nu}({\bm{x}})$.
Recall Eq. (8) that $\frac{\partial^{2}}{\partial\nu\partial{\bm{x}}}\ell_{\nu}({\bm{x}})|_{\nu=0}=\nabla\varphi({\bm{x}})$. Then Lemma 3.2, in conjunction with Lemma 3.1, indicates that the $p$th-order finite difference used in F2SA-$p$ guarantees an ${\mathcal{O}}(\nu^{p})$-approximation error to $\nabla\varphi({\bm{x}})$, which always improves the ${\mathcal{O}}(\nu)$-error guarantee of F2SA [28, 10] for any $p\geq 2$. This improved error guarantee means that we can set $\nu={\mathcal{O}}(\epsilon^{1/p})$ to obtain an ${\mathcal{O}}(\epsilon)$-accurate hyper-gradient estimator to $\nabla\varphi({\bm{x}})$, which further leads to the following improved complexity of our algorithm.
@@
For any instance in ${\mathcal{F}}^{\text{nc-sc}}(L_{0},\cdots,L_{p+1},\mu,\Delta)$ as per Definition 2.2, set the hyper-parameters as
@@
For $p=1$, our upper bound becomes $\tilde{\mathcal{O}}(\kappa^{11}\epsilon^{-6})$, which improves the $\tilde{\mathcal{O}}(\kappa^{12}\epsilon^{-6})$ bound in [10] by a factor of $\kappa$. The improvement comes from a tighter analysis in the lower-level SGD update and a careful parameter setting.
Remark 3.5 (Highly smooth region).
For $p=\Omega(\log(\nicefrac{{\kappa}}{{\epsilon}})/\log\log(\nicefrac{{\kappa}}{{\epsilon}}))$ in Definition 2.2, we can run F2SA-$q$ with $q\asymp\log(\nicefrac{{\kappa}}{{\epsilon}})/\log\log(\nicefrac{{\kappa}}{{\epsilon}})$ and the ${\mathcal{O}}(q\kappa^{9}\epsilon^{-4}(\nicefrac{{\kappa}}{{\epsilon}})^{2/q}\log(\nicefrac{{\kappa}}{{\epsilon}}))$ complexity in Theorem 3.1 simplifies to ${\mathcal{O}}(\kappa^{9}\epsilon^{-4}\log^{3}(\nicefrac{{\kappa}}{{\epsilon}})/\log\log(\nicefrac{{\kappa}}{{\epsilon}}))=\tilde{\mathcal{O}}(\kappa^{9}\epsilon^{-4})$, which matches the best-known complexity for HVP-based methods [21] under stochastic Hessian assumption (5).
In the upcoming section, we will derive an $\Omega(\epsilon^{-4})$ lower bound to prove that the F2SA-$p$ is near-optimal in the above highly smooth region if the condition number $\kappa$ is constant. We leave the study of optimal complexity for non-constant $\kappa$ to future work.
@@
Note that by Lemma 3.1 when $p$ is odd, we need to use $p+1$ points to construct the estimator, which means the algorithm needs to solve $p+1$ lower-level problems in each iteration to achieve an ${\mathcal{O}}(\nu^{p})$ error guarantee. In contrast, when $p$ is even, $p$ points are enough since the $p$th-order central difference estimator satisfies that $\alpha_{0}=0$. It suggests that even when $p$ is odd, the algorithm designed for odd $p$ may still be better. For instance, the F2SA-2 may always be a better choice than F2SA since its benefits almost come for free: (1) it still only needs to solve 2 lower-level problems as the F2SA method, which means the per-iteration complexity remains the same. (2) Although the improved complexity of F2SA-2 relies on the second-order smooth condition, without such a condition, its error guarantee in hyper-gradient estimation only degenerates to a first-order one, which means it is at least as good as F2SA.

### v2

@@ 3 The F2SA-$p$ Method
To introduce our method, we first recall the prior F2SA method (Kwon et al., 2023) and establish their relationship between finite difference schemes, which further motivates us to design better algorithms by using better finite difference formulas.
@@ 3.1 Hyper-Gradient Approximation via Finite Difference
The core idea of F2SA (Kwon et al., 2023) is to solve the reformulated penalty problem (3) and use the gradient of the penalty function to approximate the true hyper-gradient. To make connections of F2SA to the finite difference method, let us introduce the extra notation $g_{\nu}$ as the perturbed lower-lever problem with ${\bm{y}}_{\nu}^{*}({\bm{x}})$ and $\ell_{\nu}({\bm{x}})$ being its optimal solution and optimal value, respectively:
@@
Then we have $\frac{\partial}{\partial\nu}\ell_{\nu}({\bm{x}})|_{\nu=0}=\lim_{\nu\rightarrow 0}\frac{\ell_{\nu}({\bm{x}})-\ell_{0}({\bm{x}}))}{\nu}=\lim_{\nu\rightarrow 0}f({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}}))+\frac{g({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}}))-g({\bm{x}},{\bm{y}}^{*}({\bm{x}}))}{\nu}$. In our notation, we can rewrite (Chen et al., 2025b, Lemma B.3) as $\frac{\partial}{\partial\nu}\ell_{\nu}({\bm{x}})|_{\nu=0}=\varphi({\bm{x}})$. Similarly, we can also rewrite (Kwon et al., 2023, Lemma 3.1) as
@@
Let $\nu=1/\lambda$ in Eq. (3). Then the fully first-order hyper-gradient estimator (Kwon et al., 2023; Chen et al., 2025b) is exactly using forward difference to approximate $\nabla\varphi({\bm{x}})$, that is,
@@
However, the forward difference is not the only way to approximate a derivative. Essentially, it falls into a general class of $p$th-order finite difference (Atkinson and Han, 2005) that can guarantee an ${\mathcal{O}}(\nu^{p})$ approximation error. We restate this known result (with generalization to vector-valued functions) in the following lemma and provide a self-contained proof in Appendix B for completeness.
@@
Assume the function $\psi:{\mathbb{R}}\rightarrow{\mathbb{R}}^{d}$ has $C$-Lipschitz continuous $p$th-order derivative. There exist coefficients $\{\alpha_{j}\}$ such that
$\left\|\frac{1}{\nu}\sum_{j}\alpha_{j}\psi(j\nu)-\psi^{\prime}(0)\right\|={\mathcal{O}}(C\nu^{p}).$
If $p$ is even, the indices run $j=-p/2,\cdots,p/2$. If $p$ is odd, they run $j=-(p-1)/2,\cdots,(p+1)/2$. Furthermore, all the coefficients satisfy $|j\alpha_{j}|\leq 1$ for all $j\neq 0$ and $|\alpha_{0}|\leq\mathbb{I}[p\text{ is odd}]$.
The explicit formulas for $\alpha_{j}$ can be found in Appendix B. When $p=1$, we have $\alpha_{0}=-1$, $\alpha_{1}=1$, and we obtain the forward difference estimator $\psi(\nu)-\psi(0)/\nu$; When $p=2$ we have $\alpha_{-1}=-1/2,\alpha_{1}=1/2$ and we obtain the central difference estimator $(\psi(\nu)-\psi(-\nu))/(2\nu)$. Lemma 3.1 tells us that in general we can always construct a finite difference estimator ${\mathcal{O}}(\nu^{p})$ error with $p$ points for even $p$ or $p+1$ points for odd $p$ under the given smoothness conditions. Inspired by Lemma 3.1 and Eq. (8) that $\frac{\partial^{2}}{\partial\nu\partial{\bm{x}}}\ell_{\nu}({\bm{x}})|_{\nu=0}=\nabla\varphi({\bm{x}})$, we propose a fully first-order estimator via a linear combination of $\frac{\partial}{\partial{\bm{x}}}\ell_{j\nu}({\bm{x}})$ to achieve ${\mathcal{O}}(\nu^{p})$ approximation error to $\nabla\varphi({\bm{x}})$ given that $\frac{\partial^{p+1}}{\partial\nu^{p}\partial{\bm{x}}}\ell_{\nu}({\bm{x}})$ is Lipschitz continuous in $\nu$. It further leads to Algorithm 1 that will be formally introduced in the next subsection.
@@ 3.2 The Proposed Algorithm
3:  parallel for $j=-p/2,-p/2+1,\cdots,p/2$
4:   ${\bm{y}}_{t}^{j,0}={\bm{y}}_{t}^{j}$
5:   for $k=0,1,\cdots,K-1$
6:    Sample random i.i.d indexes $\{(\xi^{y}_{j},\zeta^{y}_{j})\}$.
@@
10:  end parallel for
11:  Sample random i.i.d indexes $\{({\xi_{i}^{x}},{\zeta_{i}^{x}})\}_{i=1}^{S}.$
12:  Let $\{\alpha_{j}\}_{j=-p/2}^{p/2}$ be the $p$th-order finite difference coefficients defined in Lemma 3.1.
13:  $\Phi_{t}=\frac{1}{S}\sum_{i=1}^{S}\sum_{j=-p/2}^{p/2}\alpha_{j}\left(jF_{x}({\bm{x}}_{t},{\bm{y}}_{t+1}^{j};\xi^{x}_{i})+\dfrac{G_{x}({\bm{x}}_{t},{\bm{y}}_{t+1}^{j};\zeta^{x}_{i})}{\nu}\right)$
14:  ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta_{x}{\Phi_{t}}/{\|\Phi_{t}\|}$
15:end for
Due to space limitations, we only present Algorithm 1 designed for even $p$ in the main text. The algorithm for odd $p$ can be designed similarly, and we defer the concrete algorithm to Appendix D.
Algorithm 1 follows the double-loop structure of F2SA (Chen et al., 2025b; Kwon et al., 2024a) and changes the hyper-gradient estimator to the one introduced in the previous section. Now, we give a more detailed introduction to the procedures of the two loops of F2SA-$p$.
@@
Compared to (Chen et al., 2025b; Kwon et al., 2023), the only modification we make to the outer loop is to change the gradient step to a normalized gradient step. The normalization can control the change of ${\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})$ and make the analysis of inner loops easier. We believe that all our theoretical guarantees also hold for the standard gradient step via a more involved analysis.
@@ 3.3 Complexity Analysis
This section contains the complexity analysis of Algorithm 1. We first derive the following lemma from the high-dimensional Faà di Bruno formula (Licht, 2024).
@@
Let $\nu\in(0,1/(2\kappa)]$. For any instance in the $p$th-order smooth bilevel problem class ${\mathcal{F}}^{\text{nc-sc}}(L_{0},\cdots,L_{p+1},\mu,\Delta)$ as Definition 2.2, $\frac{\partial^{p+1}}{\partial\nu^{p}\partial{\bm{x}}}\ell_{\nu}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{2p+1}\bar{L})$-Lipschitz continuous in $\nu$.
Our result generalizes the prior result for $p=1$ (Kwon et al., 2023) to any $p\in{\mathbb{N}}_{+}$ and also tightens the prior bounds for $p=2$ (Chen et al., 2025b) as we remark in the following.
@@
Note that the variables $\bm{x}$} and $\nu$ play equal roles in our analysis. Therefore, our result in $p=2$ essentially implies that $\frac{\partial^{3}}{\partial\nu\partial{\bm{x}}^{2}}\ell_{\nu}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{5}\bar{L})$-Lipschitz continuous in $\nu$ around zero, which tightens the ${\mathcal{O}}(\kappa^{6}\bar{L})$ bound of Hessian convergence in (Chen et al., 2025b, Lemma 5.1a) and is of independent interest. The main insight is to avoid the direct calculation of $\nabla^{2}\varphi({\bm{x}})=\frac{\partial^{3}}{\partial\nu\partial{\bm{x}}^{2}}\ell_{\nu}({\bm{x}})|_{\nu=0}$ which involves third-order derivatives and makes the analysis more complex, but instead always to analyze it through the limiting point $\lim_{\nu\rightarrow 0}\frac{\partial^{3}}{\partial\nu\partial{\bm{x}}^{2}}\ell_{\nu}({\bm{x}})$.
Recall Eq. (8) that $\frac{\partial^{2}}{\partial\nu\partial{\bm{x}}}\ell_{\nu}({\bm{x}})|_{\nu=0}=\nabla\varphi({\bm{x}})$. Then Lemma 3.2, in conjunction with Lemma 3.1, indicates that the $p$th-order finite difference used in F2SA-$p$ guarantees an ${\mathcal{O}}(\nu^{p})$-approximation error to $\nabla\varphi({\bm{x}})$, which always improves the ${\mathcal{O}}(\nu)$-error guarantee of F2SA (Kwon et al., 2023; Chen et al., 2025b) for any $p\geq 2$. This improved error guarantee means that we can set $\nu={\mathcal{O}}(\epsilon^{1/p})$ to obtain an ${\mathcal{O}}(\epsilon)$-accurate hyper-gradient estimator to $\nabla\varphi({\bm{x}})$, which further leads to the following improved complexity of our algorithm.
@@
For any instance in the $p$th-order smooth bilevel problem class ${\mathcal{F}}^{\text{nc-sc}}(L_{0},\cdots,L_{p+1},\mu,\Delta)$ as per Definition 2.2, set the hyper-parameters as
@@
For $p=1$, our upper bound becomes $\tilde{\mathcal{O}}(\kappa^{11}\epsilon^{-6})$, which improves the $\tilde{\mathcal{O}}(\kappa^{12}\epsilon^{-6})$ bound in (Chen et al., 2025b) by a factor of $\kappa$. The improvement comes from a tighter analysis in the lower-level SGD update and a careful parameter setting.
Remark 3.4 (Highly-smooth region).
For $p=\Omega(\log(\nicefrac{{\kappa}}{{\epsilon}})/\log\log(\nicefrac{{\kappa}}{{\epsilon}}))$ in Definition 2.2, we can run F2SA-$q$ with $q\asymp\log(\nicefrac{{\kappa}}{{\epsilon}})/\log\log(\nicefrac{{\kappa}}{{\epsilon}})$ and the ${\mathcal{O}}(q\kappa^{9}\epsilon^{-4}(\nicefrac{{\kappa}}{{\epsilon}})^{2/q}\log(\nicefrac{{\kappa}}{{\epsilon}}))$ complexity in Theorem 3.1 simplifies to ${\mathcal{O}}(\kappa^{9}\epsilon^{-4}\log^{3}(\nicefrac{{\kappa}}{{\epsilon}})/\log\log(\nicefrac{{\kappa}}{{\epsilon}}))=\tilde{\mathcal{O}}(\kappa^{9}\epsilon^{-4})$, which matches the best-known complexity for HVP-based methods (Ji et al., 2021) under stochastic Hessian assumption (5).
In the upcoming section, we will derive an $\Omega(\epsilon^{-4})$ lower bound to prove that the F2SA-$p$ is near-optimal in the above highly-smooth region if the condition number $\kappa$ is a constant. We leave the study of optimal complexity for non-constant $\kappa$ to future work.
@@
Note that by Lemma 3.1 when $p$ is odd, we need to use $p+1$ points to construct the estimator, which means the algorithm needs to solve $p+1$ lower-level problems in each iteration to achieve an ${\mathcal{O}}(\nu^{p})$ error guarantee. In contrast, when $p$ is even, $p$ points are enough since the $p$th-order central difference estimator satisfies that $\alpha_{0}=0$. It suggests that even when $p$ is odd, the algorithm designed for odd $p$ may still be better. For instance, the F2SA-2 may always be a better choice than F2SA since its benefits almost come for free: (1) it still only needs to solve 2 lower-level problems as the F2SA method, which means the per-iteration complexity remains the same. (2) Although the improved complexity of F2SA-2 relies on the second-order smooth condition, without such a condition, its error guarantee in hyper-gradient estimation will only degenerate to a first-order one, which means it is at least as good as F2SA.